Bill & Jose's Discount Cinema sold twice as many adults' tickets for $15.00 each and children's tickets for $10.00 each. Last Tuesday, a total of $1600.00 was collected from movie watchers. How many of each type of ticket were sold? (Let x represent the number of adults' tickets and y represent the number of children's tickets.)

Answer by mananth(16949)

adult tickets = x
child tickets = y
x=2y......................1
adult ticket $15
child ticket = $10
15x+10y=1600
plug value of x
15*2y+10y=1600
30y+10y=1600
40y=1600
/40
y = 40 number of child tickets
2*40 = 80 adult tickets
...
m.ananth@hotmail.ca